A 50.0kg disk accelerates from rest to an angular speed of 5.00 rad/s in 30.0 seconds. The disk's diameter is 20.0 meters.

a. Calculate the disk's moment of inertia in kgm^2.

b. Calculate the disk's angular acceleration in rad/s^2

c. Calculate the net torque mN required to accelerate the disk.

Please show all steps
Answer #1

Moment of inertia of disk=0.5 m r^2=0.5 ×50× 10^2= 2500 kg m^2

Angular acceleration=5/30=1/6 rad/ sec ^2

Torque=1/6×2500=416.66